# Learning Macroeconomic Policies based on Microfoundations: A Dynamic Stackelberg Mean Field Game Approach

### v2

@@ 5.3 Effectiveness Analysis in Diverse Households
The above experiments have already demonstrated the effectiveness of our method in addressing macroeconomic policy-making problems. However, when extending SMFG approach to real-world scenarios, where governments and households implement leader and follower policies trained via the SMFG method (referred to as SMFG government and SMFG households), there remain partial households that might not adopt the SMFG follower policy. In economic terms, this scenario can be interpreted as the government introducing new policies and providing recommendations to households, some of whom may choose not to follow these recommendations. When faced with heterogeneous households that adopt varying strategies, the question arises: is our SMFG method still effective? In experiments, to model households that do not adopt the SMFG policy, referred to as Real-data Households, we utilize behavior cloning policy trained by real data. Related experimental details are shown in Appendix D.4. We set different proportions of SMFG households (0%, 25%, 50%, 75%, and 100%) and evaluate our SMFG method using microeconomic and macroeconomic indicators. In Figure 3, each subplot displays average values for two types of households (bars on the left Y-axis) and their aggregate values (dots on the right Y-axis). GDP, a macroeconomic indicator, is shown only in aggregate, not for individual households.
Result Analysis
From Figure 3, we identify several key findings: (1) SMFG Households outperform Real-data Households across microeconomic indicators such as wealth, income, and utility, which should encourage more households to adopt the SMFG policy; (2) The presence of heterogeneous households does not prevent SMFG Households from achieving near-optimal utility, capped at 1 in the TaxAI environment; (3) The lower labor contribution from Real-data Households leads to a decrease in Per Capita GDP as their proportion increases. Thus, the more households adopt the SMFG follower policy, the more effective the SMFG method becomes.
@@ 6 Conclusion
This paper innovatively models macroeconomic policymaking as a dynamic Stackelberg mean field game, with the government acting as the leader and large-scale households as followers. We introduce a data-driven Stackelberg mean field reinforcement learning algorithm that effectively solves SMFGs. When applied to macroeconomic policy formulation, our method significantly outperforms existing economic approaches and AI-based policies, especially in scenarios involving large numbers of households. Furthermore, we demonstrate the robustness of our approach, showing that it retains effectiveness even when some households do not adopt the SMFG policy. In conclusion, this paper contributes an effective approach to modeling and solving optimal macroeconomic policies in the field of AI for economics and AI for social good.
@@ Appendix A SMFRL algorithm pseudocode
Algorithm 1 Stackelberg Mean Field Reinforcement Learning (SMFRL)
  Initialize ${Q}_{\phi^{l}},{Q}_{\phi^{l}_{-}},{Q}_{\phi^{f}},{Q}_{\phi^{f}_{-}},\pi_{\theta^{l}},\pi_{\theta^{l}_{-}},\pi_{\theta^{f}},\pi_{\theta^{f}_{-}}$.
  for epoch = 1 to M do
     Receive initial state $\mathbf{s}_{t}=\{s^{l}_{t},s^{f1}_{t},...,s^{fN}_{t}\}$.
     for  $t=1$ to max-epoch-length do
        For the leader agent, select action $a^{l}_{t}=\pi_{\theta^{l}}(s^{l}_{t})+\mathcal{N}_{t}$; for each follower agent $i$, select action $a^{fi}=\pi_{\theta^{f}}(s^{fi}_{t},a^{l}_{t})+\mathcal{N}_{t}$ w.r.t. the current policy and exploration.
        Execute actions $\mathbf{a}_{t}=\{a^{l}_{t},a^{f1}_{t},...,a^{fN}_{t}\}$ and observe reward $\mathbf{r}_{t}=\{r_{t}^{l},r^{f1}_{t},...,r^{fN}_{t}\}$ and next state $\mathbf{s}_{t+1}$.
        Store $(\mathbf{s}_{t},\mathbf{s}_{t+1},\mathbf{a}_{t},\mathbf{r}_{t},L^{f}_{t})$ in the replay buffer $\mathcal{D}$.
        $\mathbf{s}_{t}\leftarrow\mathbf{s}_{t+1}$.
        for j =1 to update-cycles do
           Sample a random minibatch of $N_{b}$ samples from $\mathcal{D}$.
           For each follower agent, perform inner updates:
           for k = 1 to inner-update-cycles do
              Set
$y^{f}_{t}=r^{f}_{t}+\gamma{Q}_{\phi^{f}_{-}}(s^{f}_{t+1},a^{f}_{t+1},a^{l}_{t+1},L^{f}_{t+1}).$
              Update follower’s critic by minimizing the loss:
$\nabla_{\phi^{f}}\mathcal{L}(\phi^{f})=\mathbb{E}_{\mathcal{D}}\Big{[}\left(y^{f}_{t}-{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{f}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{t},a^{l}_{t},L^{f}_{t})\Big{]}.$
              Update follower’s actor by policy gradient:
$\nabla_{\theta^{f}}J\approx\mathbb{E}_{\mathbf{s}_{t},a^{l}_{t},L^{f}_{t}\sim\mathcal{D}}\Big{[}\nabla_{\theta^{f}}\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})\nabla_{a^{f}_{-}}{Q}_{\phi^{f}}(s^{f}_{t},a^{f}_{-},a^{l}_{t},L^{f})|_{a^{f}_{-}=\pi_{\theta^{f}}(s^{f}_{t},a^{l}_{t})}\Big{]}.$
           end for
           After follower updates, update the leader’s networks:
           Set
$y^{l}_{t}=r_{t}^{l}+\gamma{Q}_{\phi^{l}_{-}}(s^{l}_{t+1},a^{l}_{t+1},L^{f}_{t+1})|_{a^{l}_{t+1}=\pi_{\theta^{l}_{-}}(s^{l}_{t+1})}.$
           Update leader’s critic net by minimizing the loss:
$\nabla_{\phi^{l}}\mathcal{L}(\phi^{l})=\mathbb{E}_{s^{l}_{t},a^{l}_{t},L^{f}_{t},s^{l}_{t+1}\sim\mathcal{D}}\left[\left(y^{l}_{t}-{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right)\nabla_{\phi^{l}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{t},L^{f}_{t})\right]$
           Update leader’s actor by sampled policy gradient:
$\nabla_{\theta^{l}}J\approx\mathbb{E}_{\mathbf{s}_{t}\sim\mathcal{D}}\left[\nabla_{\theta^{l}}\pi_{\theta^{l}}(s^{l}_{t})\nabla_{a^{l}_{-}}{Q}_{\phi^{l}}(s^{l}_{t},a^{l}_{-},L^{f}_{t})|_{a^{l}_{-}=\pi_{\theta^{l}}(s^{l}_{t})}\right].$
        end for
        Update target network parameters:
$\begin{gathered}\phi_{-}^{l}\leftarrow\tau_{\phi}\phi^{l}+\left(1-\tau_{\phi}\right)\phi_{-}^{l}\\ \theta_{-}^{l}\leftarrow\tau_{\theta}\theta^{l}+\left(1-\tau_{\theta}\right)\theta_{-}^{l}\\ \phi_{-}^{f}\leftarrow\tau_{\phi}\phi^{f}+\left(1-\tau_{\phi}\right)\phi_{-}^{f}\\ \theta_{-}^{f}\leftarrow\tau_{\theta}\theta^{f}+\left(1-\tau_{\theta}\right)\theta_{-}^{f}\end{gathered}$
     end for
  end for
@@ Appendix B Assumptions and Limitations
Appendix C Impact Statements
This paper presents work whose goal is to advance the fields of AI for social science and AI for economics. Our work aims to offer suggestions and references for governments and the people, yet it must not be rashly applied to the real world. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix D Additional Results
Compute Resources
Table 3: the approximate training times for several algorithms for solving dynamic SMFGs.
Algorithm N=100 (hours) N=1000 (hours)
SMFRL 4 14
MADDPG 3 16
MF-MARL 4 9
IDDPG 2 6
DDPG+Real 1 5
Rule-based+Real 0.25 2
D.1 Training Curves for Dynamic SMFGs
Due to the challenging nature of the task, the baselines we compared did not perform well in solving dynamic SMFGs. Therefore, we utilized behavior cloning as a pre-training strategy for the follower agents. Therefore, we compared SMFRL with 5 baselines in 4 different experimental setups: without behavior cloning as pre-training for follower agents at N=100 and N=1000 (marked as N=100 without BC and N=1000 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC). Figure 4 illustrates the training curves of 4 key macroeconomic indicators under these three settings. The solid line represents the average value of the metrics across the 5 random seeds, while the shaded area represents the standard deviation. Each row corresponds to one setting, and each column to a macroeconomic indicator, including per capita GDP, social welfare, income Gini, and wealth Gini. A rise in per capita GDP indicates economic growth, an increase in social welfare implies happier households and a lower Gini index indicates a fairer society. Each subplot’s Y-axis represents the indicators’ values, and the X-axis represents the training steps. Table 4 displays the test results of the 6 algorithms across 5 indicators, with each column corresponding to an experimental setting.
Figure 4 and Table 4 present two experimental findings: (1) Using BC as a pre-training method for the follower’s policy enhances the algorithms’ stability and performance. Comparing settings with and without BC (the first two rows), our method, SMFRL, shows similar convergence outcomes; however, the performance of other algorithms significantly improves across all four indicators with BC-based pre-training. Furthermore, the training curves of each algorithm are more stable. (2) The SMFRL algorithm substantially outperforms other algorithms in solving SMFGs, both in large-scale followers and without pre-training scenarios. In the setting of N=100-BC, SMFRL achieved a significantly higher per capita GDP compared to other algorithms, while its social welfare and Gini index are similar to others, essentially reaching the upper limit. Besides, in N=100 without BC and N=1000-BC, SMFRL consistently obtains the most optimal solutions across all indicators.
Table 4: Performance metrics of six algorithms across different settings on key economic indicators.
Settings Algorithm GDP per Capita Social Welfare Income Gini Wealth Gini Years
N=100 without BC Rule-based+Random 1.41e+05 69.27 0.89 0.92 1.00
DDPG+Random 1.41e+05 70.91 0.88 0.92 1.00
MADDPG 4.93e+04 55.74 0.91 0.92 1.00
IDDPG 1.21e+05 83.09 0.88 0.92 1.00
MF-MARL 8.66e+07 82.02 0.82 0.83 75.75
SMFRL 9.59e+12 96.87 0.52 0.51 300.00
N=1000 without BC Rule-based+Random 1.42e+05 705.85 0.90 0.93 1.00
DDPG+Random 1.42e+05 654.37 0.90 0.93 1.00
MADDPG 3.37e+09 645.38 0.83 0.85 63.58
IDDPG 1.33e+05 874.53 0.89 0.92 1.00
MF-MARL 1.31e+05 834.93 0.88 0.92 1.50
SMFRL 1.10e+13 968.94 0.54 0.53 300.00
N=100 with BC Rule-based+Real 3.66e+11 79.23 0.52 0.53 217.45
DDPG+Real 2.03e+12 94.50 0.46 0.48 299.85
MADDPG 6.38e+12 93.89 0.57 0.58 268.53
IDDPG 7.41e+12 98.16 0.53 0.55 300.00
MF-MARL 5.44e+12 98.21 0.50 0.52 300.00
SMFRL 1.01e+13 96.90 0.51 0.53 299.89
N=1000 with BC Rule-based+Real 1.41e+05 334.79 0.90 0.93 1.00
DDPG+Real 4.92e+11 527.09 0.75 0.79 100.68
MADDPG 6.82e+12 954.88 0.56 0.62 278.50
IDDPG 2.79e+12 512.19 0.77 0.81 100.68
MF-MARL 1.13e+05 440.00 0.90 0.93 1.00
SMFRL 9.68e+12 975.15 0.52 0.51 300.00
D.2 Training Curves for Various Tax Policies
We compare the performance of 6 policies across four economic indicators under two settings: with N=100 and N=1000 households. Figure 5 displays the training curves and Table 2 shows the test results in Appendix. Both Figure 5 and Table 2 indicate that the SMFG method significantly surpasses other policies in the task of optimizing GDP, and achieves the highest social welfare. When N=100, the Saez tax achieves the lowest income and wealth Gini coefficients, suggesting greater fairness. However, at N=1000, SMFG performs optimally across all economic indicators, while the effectiveness of other policies noticeably diminishes as the number of households increases. The Saez tax also reduces the Gini index, but not as effectively or stably as the SMFG.
D.3 Efficiency-Equity Tradeoff of Policies
D.4 Behavior Cloning Experiments
We conduct behavior cloning based on real data to simulate the behavior strategies of households in realistic scenarios, which are then used in Experiment 5.3 to compare with SMFG Households. We use the statistical data from the 2022 Survey of Consumer Finances (SCF)  [^3] https://www.federalreserve.gov/econres/scfindex.htm. Details about the behavior cloning experiments are provided below. This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), MF-MARL, and MADDPG network. The first two, as their network outputs, are probability distributions, use negative log-likelihood loss (Figure 7 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 7 right). The loss convergence curve of behavior cloning is shown in Figure 7. It can be observed that the AI economist’s network, due to its complexity, struggles to converge to near -1 like MLP. The losses corresponding to MFRL and MADDPG can converge to below 0.1.
Figure 7: The behavior cloning loss for 4 networks in two loss types.
Appendix E Saez tax
@@
The Saex tax policy is often considered a suggestion for specific tax reforms in the real world. The specific calculation method is as follows [50]. The Saez tax utilizes income distribution $f(z)$ and cumulative distribution $F(z)$ to get the tax rates. The marginal tax rates denoted as $\tau(z)$, are expressed as a function of pretax income $z$, incorporating elements such as the income-dependent social welfare weight $G(z)$ and the local Pareto parameter $\alpha(z)$.
@@
Appendix F Theoretical Proof
Proposition F.1.
The action value functions $Q_{i}$ and ${Q}_{t}$ are equivalent, i.e. $Q_{i}(\mathbf{s}_{t},\mathbf{a}_{t})={Q}_{t}(s_{t}^{l},a_{t}^{l},L_{t}^{f})$, $\forall\mathbf{s}_{t}\in\mathcal{S}^{l}\times\{\mathcal{S}^{f}\}^{N}$, $\forall\mathbf{a}_{t}\in\mathcal{A}^{l}\times\{\mathcal{A}^{f}\}^{N}$, $\forall L_{t}^{f}\in\mathbb{P}(\mathcal{S}^{f}\times\mathcal{A}^{f})$.
Proof.
We commence by defining a partial order relation $\prec$ on $\mathcal{S}^{f}\times\mathcal{A}^{f}$. For any given vector $\mathbf{v}$, we construct a sorted counterpart $\tilde{\mathbf{v}}$, which maintains the same elements as $\mathbf{v}$ but orders them in ascending sequence based on $\prec$. For example, if $\mathbf{v}=\{(s_{1},a_{1}),(s_{2},a_{2}),(s_{3},a_{3})\}$, then $\tilde{\mathbf{v}}$ is arranged as $\{(s_{(1)},a_{(1)}),(s_{(2)},a_{(2)}),(s_{(3)},a_{(3)})\}$, where $(s_{(1)},a_{(1)})\prec(s_{(2)},a_{(2)})\prec(s_{(3)},a_{(3)})$). Exploiting the homogeneity of the follower agents, we can equate $Q(\mathbf{s}_{t},\mathbf{a}_{t})$ with $Q(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})$, where $\mathbf{v}$ encapsulates the joint state-action pairs of the followers, and $\tilde{\mathbf{v}}$ represents its sorted form. We further define $\mathcal{\mathbf{v}}$ as the collection of all such sorted vectors, and $\mathcal{L}$ as the ensemble of mean-field distributions. The empirical distribution of the followers’ state-action pairs, denoted as $L:\mathcal{V}\rightarrow\Delta(\mathcal{V})$, is expressed as $L^{f}_{t}(s,a)=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}_{\{(\tilde{s}_{i},\tilde{a}_{i})=(\tilde{s},\tilde{a})\}}$. The bijectiveness of the mapping $L$, ensured by the sorting process, leads to the conclusion that $Q_{t}(\mathbf{s}_{t},\mathbf{a}_{t})=Q_{t}(s^{l}_{t},a^{l}_{t},\mathbf{v})=Q_{t}(s^{l}_{t},a^{l}_{t},\tilde{\mathbf{v}})={Q}_{t}(s^{l}_{t},a^{l}_{t},L^{f}_{t})$. ∎
Appendix G Economic Model Details

### v3

@@ 5 Experiment
Ablation Studies of SMFG
We conduct ablation studies to analyze the impact of the leader-follower update and mean field approximation modules in the SMFG method on macroeconomic indicators, as well as on convergence and social welfare.
Baselines
We conduct comparisons based on the SMFG method by excluding the leader-follower update (denoted as S) and the mean field approximation (denoted as MF) individually:
1. SMFG-S: SMFG without leader-follower update. The government agent uses the independent DDPG (de Witt et al. 2020), while interactions among large-scale households are modeled as a mean field game and trained using the mean-field MARL algorithm (Yang et al. 2018).
2. SMFG-MF: SMFG without mean field approximation. Both the government and household agents are trained using a variant of MADDPG algorithm (Lowe et al. 2017), with the government making decisions first, followed by households.
3. SMFG-S-MF: SMFG without both leader-follower update and mean field approximation. In this case, both the government and household agents use independent DDPG.
Macroeconomic Indicators Analysis
Figure 4 presents the results of the ablation studies on various macroeconomic indicators, evaluated across five seeds. The height of each bar represents the mean of the macroeconomic indicators, while the error bars indicate the variance. Our findings reveal that: (1) Excluding either module leads to a significant performance drop compared to the full SMFG method, with this effect becoming more pronounced as the number of agents $N$ increases. This underscores the critical importance of both the S and MF modules. (2) The SMFG-S variant outperforms SMFG-MF and SMFG-S-MF in terms of higher Per Capita GDP, social welfare, and Years, as well as lower Gini index. This suggests the dimensionality reduction of the MF module is crucial for the SMFG method’s overall effectiveness. (3) The low variance of the SMFG method indicates its stability.
Convergence and Social Welfare
Table 2 presents the performance of different baselines across three game theory-related metrics: leader’s payoff, social welfare, and exploitability. Leader’s payoff reflects the performance of the leader agent (i.e., the government), social welfare corresponds to the total utility of all households, and exploitability measures the distance between current policy and Nash equilibrium. The methods for calculating the metrics and optimal values are provided in Appendix D.
From Table 2, we find: (1) SMFG achieves the highest leader’s payoff, the highest social welfare, and near-zero exploitability, indicating that the SMFG policy is closest to both the equilibrium and the maximal social welfare. (2) The high exploitability exhibited by SMFG-S indicates that the leader-follower update module facilitates the algorithm’s convergence toward equilibrium. (3) SMFG-MF outperforms SMFG-S-MF in terms of the leader’s payoff, suggesting that the leader-follower update module aids in optimizing the leader’s objective. However, it is challenging for SMFG-MF to optimize the followers’ social welfare.
Leader’s Exploitability Social
Payoff Welfare
[…]
Optimal Value \ \ 0. 0. 100 1000
SMFG-S-MF -774 -716 0.652 1.023 83 859
SMFG-S -448 -856 3.161 1.213 82 782
SMFG-MF -535 -441 0.782 0.725 54 499
SMFG (ours) 3294 3376 0.002 0.023 98 971
Table 2: Ablation studies of the SMFG method based on game theory metrics for N=100 and N=1000. Optimal values are provided for reference.
Effectiveness Analysis of SMFG
@@
The above experiments have shown the effectiveness of the SMFG method in macroeconomic policymaking. Assuming the government is willing to implement macroeconomic policies learned by the SMFG method, we aim to explore the impact of non-adoption by some households. In our experiments, we introduce Non-SMFG Households whose policies are trained using a behavior cloning algorithm based on real data, reflecting the decision-making styles of real-world households. Details of these behavior cloning experiments are detailed in Appendix E. Additionally, we denote governments and households that adopt the SMFG policy as SMFG Government and SMFG Households, respectively. We set different proportions of SMFG households (0%, 25%, 50%, 75%, and 100%) and evaluate our SMFG method using microeconomic and macroeconomic indicators. The results is shown in Figure 5, each subplot displays average values for two types of households (bars on the left Y-axis) and their aggregate values (dots on the right Y-axis). GDP, a macroeconomic indicator, is shown only in aggregate, not for individual households.
Results
From Figure 5, we observe several intriguing findings: (1) SMFG Households outperform Non-SMFG Households across microeconomic indicators such as wealth, income, and utility, suggesting that the policy of SMFG households is more attractive and likely to encourage broader adoption; (2) The presence of Non-SMFG Households does not prevent SMFG Households from achieving near-optimal utility, which is capped at 1 in the TaxAI environment; (3) As the proportion of SMFG Households increases, the per capita GDP also rises, indicating that the government should encourage households to adopt the SMFG policy.
@@ 6 Conclusion
Addressing issues in macroeconomic policymaking, this paper innovatively proposes the SMFG method, which includes dynamic SMFG for modeling and SMFRL algorithm for solving. The SMFG method optimizes government’s policy by considering the dynamic responses of large-scale households to policy shifts, ultimately deriving optimal policies for both government and households. Experimental results demonstrate the SMFG method’s superior performance, stability, and convergence over existing approaches. When extended to real-world scenarios, even if some households do not adopt the SMFG policy, their wealth and utility are not as high as those who do, encouraging more households to adopt the SMFG policy. In conclusion, this paper contributes an effective approach to modeling and solving optimal macroeconomic policies in the fields of AI for economics and AI for social impact.
@@ Appendix B Assumptions and Limitations
Appendix C Saez tax
@@
The Saex tax policy is often considered a suggestion for specific tax reforms in the real world. The specific calculation method is as follows (Saez 2001). The Saez tax utilizes income distribution $f(z)$ and cumulative distribution $F(z)$ to get the tax rates. The marginal tax rates denoted as $\tau(z)$, are expressed as a function of pretax income $z$, incorporating elements such as the income-dependent social welfare weight $G(z)$ and the local Pareto parameter $\alpha(z)$.
@@
Appendix D Game Theory Metrics
We will utilize the following metrics related to game theory to evaluate the effectiveness of the leader and follower policies: (1) The leader’s payoff, which indicates the performance of the leader’s policy in optimizing the leader’s objective; (2) Exploitability, which measures the deviation of the agent’s policy from the best response; (3) Social welfare, which assesses the deviation of the current state from the social optimum.
Leader’s Payoff
Exploitability
Exploitability is a critical metric in evaluating the convergence of policies and quantifying the divergence from the best response strategy in game theory. For a follower, exploitability $\mathcal{E}^{f}(\pi^{f};\pi^{l})$ is defined as the difference in payoffs between the follower’s actual policy $\pi^{f}$ and its optimal response ${\pi^{f}}^{*}$, given the leader’s policy $\pi^{l}$. Formally, it is represented as:
$\mathcal{E}^{f}(\pi^{f};\pi^{l})=J^{f}(\pi^{l},{\pi^{f}}^{*},L)-J^{f}(\pi^{l},\pi^{f},L),$
where $J^{f}$ denotes the cumulative reward for the follower, defined in Section 3.
Similarly, the leader’s exploitability $\mathcal{E}^{l}(\pi^{l};\pi^{f})$ measures the payoff difference between the leader’s policy $\pi^{l}$ and its best response ${\pi^{l}}^{*}$, given the followers’ response policy $\pi^{f}$ and state-action distribution $L$. This is given by:
$\mathcal{E}^{l}(\pi^{l};\pi^{f})=J^{l}(\pi^{l*},\pi^{f},L)-J^{l}(\pi^{l},\pi^{f},L),$
with $J^{l}$ representing the cumulative reward for the leader, and $L=\{L_{t}\}_{t=0}^{T}$ detailing the state-action distribution for followers over time (see Section 3).
The overall exploitability, which measures the discrepancy from Nash equilibrium for both the leader and the followers, is defined as:
$\mathcal{E}(\pi^{l},\pi^{f})=\mathcal{E}^{f}(\pi^{f};\pi^{l})+\mathcal{E}^{l}(\pi^{l};\pi^{f}),$
A near-zero value of $\mathcal{E}(\pi^{l},\pi^{f})$ indicates that the policies of both the leader and the followers are approaching their respective optimal strategies ${\pi^{l}}^{*}$ and ${\pi^{f}}^{*}$, signifying an equilibrium state.
Social Optimum and Social Welfare
In economic theory, the Social Optimum describes a state in which the allocation of resources achieves maximum efficiency, as measured by social welfare (Arrow and Debreu 1954; Malul and Bar-El 2009). Given the leader’s policy $\pi^{l}$ and the representative follower’s policy $\pi^{f}$ among large-scale followers, social welfare $\mathcal{SW}(\pi^{l},\pi^{f})$ is approximately calculated as the sum of the utility functions defined in Section 3 of the $N$ followers:
$\mathcal{SW}(\pi^{l},\pi^{f})$ $=\sum^{N}_{i=1}J^{fi}\left(\pi^{l},\pi^{f},L\right)$
$=\mathbb{E}_{s_{0}^{f}\sim\mu_{0}^{f},\mathbf{s}_{t+1}\sim P}\left[\sum_{t=0}^{T}\sum^{N}_{i=1}r^{fi}(\mathbf{s}_{t},a^{l}_{t},a^{fi}_{t},L_{t})\right]$
Appendix E Additional Results
Compute Resources
Algorithm N=100 (hours) N=1000 (hours)
SMFG 4 14
SMFG-MF 3 16
SMFG-S 4 9
SMFG-S-MF 2 6
Free Market 0.25 2
Saez Tax 4 23
AI Economist 6.5 N/A
Table 3: The approximate training times for baselines in our experiments.
Experimental Results of Ablation Studies
In this section, we present additional experimental results from ablation studies, including training curves 6 and a test table 4, as well as experiments incorporating the use of behavior cloning as a pre-training strategy for follower agents. We find that the SMFG method without behavior cloning as pre-training still surpasses other baselines that utilize behavior cloning. More specifically, we compared SMFG with 5 baselines across 4 different experimental setups: without behavior cloning as pre-training for follower agents at N=100 and N=1000 (marked as N=100 without BC and N=1000 without BC); with BC-based pre-training for follower agents at N=100 and N=1000 (N=100-BC; N=1000-BC). Figure 6 illustrates the training curves of 4 key macroeconomic indicators under these four settings. The solid line represents the average value of the metrics across the 5 random seeds, while the shaded area represents the standard deviation. Each row corresponds to one setting, and each column to a macroeconomic indicator, including per capita GDP, social welfare, income Gini, and wealth Gini. A rise in per capita GDP indicates economic growth, an increase in social welfare implies happier households and a lower Gini index indicates a fairer society. Each subplot’s Y-axis represents the indicators’ values, and the X-axis represents the training steps. Table 4 displays the test results of the 6 algorithms across 5 indicators, with each column corresponding to an experimental setting.
Figure 6 and Table 4 present two experimental findings: (1) Using BC as a pre-training method for the follower’s policy enhances the algorithms’ stability and performance. Comparing settings with and without BC (the first two rows), our method, SMFG, shows similar convergence outcomes; however, the performance of other algorithms significantly improves across all four indicators with BC-based pre-training. Furthermore, the training curves of each algorithm are more stable. (2) The SMFG method substantially outperforms other algorithms in solving SMFGs, both in large-scale followers and without pre-training scenarios. In the setting of N=100-BC, SMFG achieved a significantly higher per capita GDP compared to other algorithms, while its social welfare and Gini index are similar to others, essentially reaching the upper limit. Besides, in N=100 without BC and N=1000-BC, SMFG consistently obtains the most optimal solutions across all indicators.
Settings Algorithm Per Capita GDP Social Welfare Income Gini Wealth Gini Years
N=100 without BC Rule-based+Random 1.41e+05 69.27 0.89 0.92 1.00
DDPG+Random 1.41e+05 70.91 0.88 0.92 1.00
SMFG-MF 1.23e+05 48.17 0.89 0.93 1.00
SMFG-S-MF 1.21e+05 83.09 0.88 0.92 1.00
SMFG-S 8.66e+07 82.02 0.82 0.83 75.75
SMFG 9.59e+12 96.87 0.52 0.51 300.00
N=1000 without BC Rule-based+Random 1.42e+05 705.85 0.90 0.93 1.00
DDPG+Random 1.42e+05 654.37 0.90 0.93 1.00
SMFG-MF 1.48e+05 499.01 0.89 0.93 1.02
SMFG-S-MF 1.33e+05 874.53 0.89 0.92 1.00
SMFG-S 1.31e+05 834.93 0.88 0.92 1.50
SMFG 1.10e+13 968.94 0.54 0.53 300.00
N=100 with BC Rule-based+Real 3.66e+11 79.23 0.52 0.53 217.45
DDPG+Real 2.03e+12 94.50 0.46 0.48 299.85
SMFG-MF 6.38e+12 93.89 0.57 0.58 268.53
SMFG-S-MF 7.41e+12 98.16 0.53 0.55 300.00
SMFG-S 5.44e+12 98.21 0.50 0.52 300.00
SMFG 1.01e+13 96.90 0.51 0.53 299.89
N=1000 with BC Rule-based+Real 1.41e+05 334.79 0.90 0.93 1.00
DDPG+Real 4.92e+11 527.09 0.75 0.79 100.68
SMFG-MF 6.82e+12 954.88 0.56 0.62 278.50
SMFG-S-MF 2.79e+12 512.19 0.77 0.81 100.68
SMFG-S 1.13e+05 440.00 0.90 0.93 1.00
SMFG 9.68e+12 975.15 0.52 0.51 300.00
Table 4: Performance metrics of six algorithms across different settings on key macroeconomic indicators.
Training Curves for Various Tax Policies
We compare the performance of 6 policies across four economic indicators under two settings: with N=100 and N=1000 households. Figure 7 displays the training curves and Table 1 shows the test results. Both Figure 7 and Table 1 indicate that the SMFG method significantly surpasses other policies in the task of optimizing GDP, and achieves the highest social welfare. When N=100, the Saez tax achieves the lowest income and wealth Gini coefficients, suggesting greater fairness. However, at N=1000, SMFG performs optimally across all economic indicators, while the effectiveness of other policies noticeably diminishes as the number of households increases. The Saez tax also reduces the Gini index, but not as effectively or stably as the SMFG.
Efficiency-Equity Tradeoff of Policies
Behavior Cloning Experiments
We conduct behavior cloning based on real data to simulate the behavior strategies of households in realistic scenarios, which are then used in Experiment 5 to compare with SMFG Households. We collect the statistical data from the 2022 Survey of Consumer Finances (SCF) (https://www.federalreserve.gov/econres/scfindex.htm) as the real data buffer $\mathcal{D}_{real}$.
Based on real data, we fetch a large number of followers’ state-action pairs $\{s^{f},a^{f}\}$ from a real-data buffer $\mathcal{D}_{real}$ for behavior cloning. For different settings of network structures, we have chosen two types of loss: when the neural network outputs a probability distribution of actions, we use the negative log-likelihood loss (NLL loss); when the neural network outputs action values, we employ the mean square error loss (MSE loss). Our goal is to find the optimal parameters $\theta$ as the follower’s policy network $\pi_{\theta}$ initialization, thereby minimizing the loss to its lowest convergence.
$\min_{\theta}\mathcal{L}_{NLL}$ $=-\mathbb{E}_{s^{f},a^{f}\sim\mathcal{D}}\log{\pi_{\theta}(a^{f}\mid s^{f})},$
$\min_{\theta}\mathcal{L}_{MSE}$ $=\mathbb{E}_{s^{f},a^{f}\sim\mathcal{D}}\left(a^{f}-a\right)^{2}|_{a=\pi_{\theta}(s^{f})}.$
This experiment conducts behavior cloning on networks for four different household policies: Multilayer Perceptron (MLP), AI economist’s network (MLP+LSTM+MLP), SMFG-S, and SMFG-MF network. The first two, as their network outputs, are probability distributions, use negative log-likelihood loss (Figure 9 left); the latter two’s networks employ deterministic policies, hence they use mean square error loss against real data (Figure 9 right). The loss convergence curve of behavior cloning is shown in Figure 9. It can be observed that the AI economist’s network, due to its complexity, struggles to converge to near -1 like MLP. The losses corresponding to MFRL and SMFG-MF can converge to below 0.1.
Figure 9: The behavior cloning loss for 4 networks in two loss types.
Appendix F Economic Model Details